# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ Abstract
Neuronal ensemble activity, including coordinated and oscillatory patterns, exhibits hallmarks of nonequilibrium systems with time-asymmetric trajectories to maintain their organization. However, assessing time asymmetry from neuronal spiking activity remains challenging. The kinetic Ising model provides a framework for studying the causal, nonequilibrium dynamics in spiking recurrent neural networks. Recent theoretical advances in this model have enabled time-asymmetry estimation from large-scale steady-state data. Yet, neuronal activity often exhibits time-varying firing rates and coupling strengths, violating the steady-state assumption. To overcome these limitations, we developed a state-space kinetic Ising model that accounts for non-stationary and nonequilibrium properties of neural systems. This approach incorporates a mean-field method for estimating time-varying entropy flow, a key measure for maintaining the system’s organization by dissipation. Applying this method to mouse visual cortex data revealed greater variability in causal couplings during task engagement despite the reduced neuronal activity with increased sparsity. Moreover, higher-performing mice exhibited increased coupling-related entropy flow per spike during task engagement, suggesting economical computation in the higher-performing mice. These findings underscore the model’s utility in uncovering intricate asymmetric causal dynamics in neuronal ensembles and linking them to behavior through the thermodynamic underpinnings of neural computation.
@@ Introduction
Neural systems are no exception. In animals engaged in behavioral and cognitive tasks, the dynamics of neuronal population activity exhibit hallmarks of nonequilibrium systems. Notable examples include the rotational activity of M1 neurons during motor execution tasks [16, 17] and the sequential patterns observed in hippocampal neurons, including their replay, during navigation and sleep [18, 19, 20]. Since the original proposal of cell assembly and its phase sequences by Donald O. Hebb [21], coordinated sequential patterns have been thought fundamental for memory consolidation and retrieval [22, 23, 24, 25]. Recently, studies on fMRI or ECoG suggested that increased time-asymmetry in neural signals, quantified by steady-state entropy production [7, 26, 27], could serve as a signature of consciousness [28, 29, 30, 31] or reflect cognitive load demanded by tasks [32]. For instance, entropy production measured from ECoG signals of non-human primates is diminished during sleep and certain types of anesthesia in comparison with wakefulness [28, 31], indicating that the awake state includes more directed temporal patterns. However, assessing entropy production directly from neuronal spiking activities remains challenging. Further complicating this issue, neural signals exhibit non-stationary dynamics, including oscillatory behavior, which hinders the use of steady-state entropy production metrics.
@@
The kinetic Ising model is a prototypical model of recurrent neural networks [33, 34]. It extends the equilibrium Ising model, which has been successfully applied to empirical spiking data to elucidate the thermodynamic and associative-memory properties of neural systems [35, 36]. In the kinetic Ising system, neurons are causally driven by the past states of self and other neurons as well as a force representing intrinsic excitability of the neurons and/or an influence of unobserved concurrent signals. When neurons receive steady inputs and their causal couplings are asymmetric, the system does not relax to an equilibrium state. Instead, it exhibits steady-state nonequilibrium dynamics characterized by non-zero entropy production. Recent theoretical studies on steady-state entropy production have elucidated its behavior in relation to distinct phases of the Ising system, including critical phase transitions [37]. Mean-field theories have been developed for kinetic Ising systems [38, 39, 40, 41, 42, 43], enabling the estimation of steady-state entropy production from large-scale spike sequences [43]. However, neuronal activity exhibits dynamical changes not only in firing rates but also in the strength of their interactions, both of which violate the steady-state assumptions.
@@
To account for the non-stationary dynamics of neural systems, the state-space method [44, 45] has been applied to the Ising system [46, 47, 48, 49, 50]. In these approaches, Bayesian filtering and smoothing algorithms have been developed to estimate time-dependent parameters of the Ising model, along with an EM algorithm [51, 52] to optimize various hyper-parameters. These models have enabled researchers to trace time-varying neuronal interactions while neurons’ internal parameters change dynamically, absorbing the effect from unobserved concurrent signals. Additionally, it has elucidated the thermodynamic quantities of neural systems (e.g., free energy and specific heat) in a time-dependent manner, in relation to the behavioral paradigms of tasks [48]. Nevertheless, these methods assume an equilibrium Ising model with symmetric couplings, which limits their ability to assess the nonequilibrium properties of observed neural activities.
@@
In this study, we develop the state-space kinetic Ising model to account for the non-stationary and nonequilibrium properties of neural activities and construct a mean-field method for estimating time-varying entropy flow, an essential component of entropy production that quantifies the dissipation of entropy, from spiking activities of neural ensembles. Given this method, we hypothesize that entropy flow, as estimated with the kinetic Ising framework, reflects the capacity of neural populations to perform meaningful computation under energetic and behavioral-time constraints [53]. Specifically, we expect that high-performing animals would exhibit greater entropy flow per spike, consistent with efficient coding.
@@
Application of the methods to mice V1 neurons revealed behavior-dependent changes in entropy flow. From the analysis of 37 mice, we found that while spike rates of the populations are lower on average and exhibited sparser distributions when mice actively engaged in tasks than in passive conditions, active engagement significantly enhanced the variability of the neuronal couplings, which contributed to increasing entropy flow. Further, higher-performing mice exhibited stronger entropy flow per spike in active engagement than in passive condition. We corroborated contributions of couplings to this tendency using trial-shuffled data that excluded influences of firing-rate dynamics and sampling errors in estimating neuronal couplings. Thus, the method enabled us to reveal contributions of behavior-related neuronal couplings to the causal activities in sparsely active neuronal populations, while isolating firing-rate dynamics. These results imply economical representation of stimuli by time-asymmetric causal activity in competent mice.
@@
This paper is organized as follows. In Results, we first introduce the state-space kinetic Ising model and its estimation method. Next, we introduce the mean-field method for estimating entropy flow. We validate these methods through simulations and then apply them to mice V1 data. Finally, we relate our findings with previous studies and discuss their implications in sparse population coding.
@@ Results
Model specifications
In neurophysiological experiments, the experimentalists simultaneously record the activity of multiple neurons while animals are exposed to a stimulus or perform a task and repeat the recordings multiple times under the same experimental conditions. We analyze the quasi-simultaneous activity of neurons using binarized spike sequences. For this goal, we convert the simultaneous sequences of spike timings of $N$ neurons into sequences of binary patterns by binning them with a bin width of $\Delta$ [ms]. We assign a value of $1$ if there is one or more spikes in a bin and 0 otherwise. We assume that there are $T+1$ bins for each trial with an initial bin being 0-th bin, and $L$ trials in total. Below, we treat the bins as discrete time steps and refer to the $t$-th bin as time $t$. We let $x^{l}_{i,t}=\{0,1\}$ be a binary variable at time $t$ in the $l$-th trial ($i=1,\ldots,N$, $t=0,\ldots,T$, $l=1,\ldots,L$). We collectively denote the binary patterns of simultaneously recorded neurons at time $t$ in the $l$-th trial using a vector, $\mathbf{x}^{l}_{t}=(x^{l}_{1,t},\ldots,x^{l}_{i,t},...x^{l}_{N,t})$. Further, we denote the patterns at time step $t$ from all trials by $\mathbf{x}_{t}=(\mathbf{x}^{1}_{t},\ldots,\mathbf{x}^{l}_{t},\ldots,\mathbf{x}^{L}_{t})$ and denote all the patterns up to time $t$ by $\mathbf{x}_{0:t}$.
@@
where $\theta_{i,t}$ is a time-dependent (external) field parameter that determines the bias for inputs to the $i$-th neuron at time $t$ and $\theta_{ij,t}$ is a time-dependent coupling parameter from the $j$-th neuron to the $i$-th neuron. These parameters are collectively denoted as $\boldsymbol{\theta}_{t}=(\boldsymbol{\theta}^{1}_{t},\ldots,\boldsymbol{\theta}^{i}_{t},\ldots,\boldsymbol{\theta}^{N}_{t})$ and $\boldsymbol{\theta}^{i}_{t}=(\theta_{i,t},\theta_{i1,t},\ldots\theta_{ij,t},\ldots{\theta}_{iN,t})$. Furthermore, $\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$ is the log normalization term defined as
@@
We also specify $p(\mathbf{x}_{0})$, a probability mass function of the binary patterns at time $t=0$, which we assume $p(\mathbf{x}_{0})=\prod_{i=1}^{N}\prod_{l=1}^{L}p(x^{l}_{i,0})$, where $p(x^{l}_{i,0})=0.5$.
@@
$p(\boldsymbol{\theta}_{1:T}|\mathbf{w})$ $=\prod_{i=1}^{N}\left[p(\boldsymbol{\theta}_{1}^{i}|\boldsymbol{\mu}^{i},\mathbf{\Sigma}^{i})\prod_{t=2}^{T}p(\boldsymbol{\theta}_{t}^{i}|\boldsymbol{\theta}_{t-1}^{i},\mathbf{Q}^{i})\right],$ (3)
where $\mathbf{w}$ denotes the collection of the hyper-parameters: $\mathbf{w}=[\boldsymbol{\mu}^{1},\ldots,\boldsymbol{\mu}^{N},\boldsymbol{\Sigma}^{1},\ldots,\boldsymbol{\Sigma}^{N},\mathbf{Q}^{1},\ldots,\mathbf{Q}^{N}]$. Namely, we assume independence of the parameters of a neuron from those of the other neurons, which significantly reduces computational costs. The transition of the $i$-th neuron follows the linear Gaussian models:
$p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1},\mathbf{Q}^{i})$ $=\frac{1}{\sqrt{|2\pi\mathbf{Q}^{i}|}}\exp\left[\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})\right],$ (4)
while the initial density is given by $p(\boldsymbol{\theta}_{1}^{i}|\boldsymbol{\mu}^{i},\mathbf{\Sigma}^{i})=\mathcal{N}(\boldsymbol{\mu^{i}},\boldsymbol{\Sigma^{i}})$.
Estimating the time-dependent parameters using optimized hyper-parameters
@@
while optimizing the parameter $\mathbf{w}$ under the principle of maximizing marginal likelihood:
@@
The Expectation-Maximization (EM) algorithm [54] offers a way to construct the approximate posterior with optimized hyperparameters by alternately constructing the approximate posterior density while fixing the parameters (E-step) and optimizing the parameters while fixing the approximate posterior (M-step). The construction of the approximate posterior density at the E-step is performed by sequentially applying Bayes algorithms in a forward and backward manner, where we approximate the posteriors by Gaussian distributions using the Laplace’s method. Thus, the method yields the mean and variance of the approximated Gaussian posterior at time $t$, which are denoted as $\boldsymbol{\theta}_{t|T}$ and $\mathbf{W}_{t|T}$, respectively. See Methods and Supporting Information S1 text for the details of the algorithm.
Estimating entropy flow
Using the learned parameters $\boldsymbol{\theta}_{1:T}$ of the kinetic Ising model from spike data, we estimate entropy flow (also known as bath entropy change) at each time step. The entropy flow at time $t$ is defined as:
$\sigma_{t}^{\rm flow}=\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}p(\mathbf{x}_{t},\mathbf{x}_{t-1})\log\frac{p(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})}{p(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})},$ (7)
where $p(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ represents the probability of observing time-reversed processes generated under the forward model. Because we use the natural logarithm, we report entropy flow in units of nats. Eq. 7 is related to the entropy production $\sigma_{t}$ at time $t$ [7, 9, 26, 27] as follows:
$\sigma_{t}$ $=\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}p(\mathbf{x}_{t},\mathbf{x}_{t-1})\log{\frac{p(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})p_{t-1}(\mathbf{x}_{t-1})}{p(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})p_{t}(\mathbf{x}_{t})}}$
@@
The entropy production is non-negative: $\sigma_{t}\geq 0$. Thus, the positive entropy flow allows a decrease in the system’s entropy: namely, the system can be more structured or organized when the entropy flow is positive. Since it is challenging to estimate the system’s entropy or its change, here we estimate the entropy flow, which provides the lower bound of the entropy change: $S_{t}-S_{t-1}\geq-\sigma_{t}^{\rm flow}$. Similarly, since the total entropy production across all time steps ${\sigma}_{1:T}$ is given as
@@
In this study, we refer to Eq. 7 as ‘entropy flow’ because it is related to heat flow to reservoirs (thermal bath) and the entropy change of the reservoirs in thermodynamics [53]. We note that Eq. 7 differs from the entropy flow defined in [55, 56], which was obtained by the decomposition of the dissipation function [27] as an alternative of entropy production. See [27, 57] for their distinct definitions and decompositions for the case of discrete-time systems.
@@
$\sigma_{t}^{\rm flow}$ $=\sum_{i}\theta_{i,t}\left(E_{\mathbf{x}_{t}}x_{i,t}-E_{\mathbf{x}_{t-1}}x_{i,t-1}\right)+\sum_{i,j}\theta_{ij,t}E_{\mathbf{x}_{t},\mathbf{x}_{t-1}}\left(x_{i,t}x_{j,t-1}-x_{i,t-1}x_{j,t}\right)$
$\phantom{===}-\sum_{i}\left(E_{\mathbf{x}_{t-1}}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{t-1})-E_{\mathbf{x}_{t}}\psi(\boldsymbol{\theta}^{i}_{t-1},\mathbf{x}_{t})\right),$ (11)
where $E_{\mathbf{x}_{t}}$ and $E_{\mathbf{x}_{t},\mathbf{x}_{t-1}}$ represents expectation by $p(\mathbf{x}_{t})$ and $p(\mathbf{x}_{t},\mathbf{x}_{t-1})$, respectively. In the steady state with time-independent parameters, only the second term remains, which can be quantified by the couplings and delayed correlations: $\sigma_{t}^{\rm flow}=(\sum_{ij}\theta_{ij}-\theta_{ji})D$, where $D_{ij,t}=E_{\mathbf{x}_{t},\mathbf{x}_{t-1}}x_{i,t}x_{j,t}-E_{\mathbf{x}_{t}}x_{i,t}E_{\mathbf{x}_{t}}x_{j,t}$ [43].
Mean-field method for estimating entropy flow
Entropy flow (Eq. 7) requires the expectation by the joint density $p(\mathbf{x}_{t},\mathbf{x}_{t-1})$, which is time-consuming for large systems. While the mean-field methods for the kinetic Ising model [38, 39, 40, 41, 42, 43] were employed to estimate steady-state entropy flow [43], the mean-field method for estimating time-varying entropy flow remains unexplored. Here we develop the mean-field method for estimating dynamic entropy flow.
The entropy flow $\sigma_{t}^{\rm flow}$ can be decomposed into the forward and reversed components,
@@
where $\sigma_{t}^{\rm forward}$ and $\sigma_{t}^{\rm backward}$ are the entropy of the conditional and time-reversed conditional distribution, respectively. The proposed mean-field method estimates the entropy flow by approximating the forward and time-reversed conditional entropy using the Gaussian integral:
@@
${\sigma}^{\rm backward}_{t}$ $\approx\sum_{i=1}^{N}\int\mathcal{D}_{z}\,\phi_{i}(g_{i,t,t}+z\sqrt{\Delta_{i,t,t}}),$ (14)
where $\mathcal{D}_{z}=\frac{\mathrm{d}z}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}z^{2}\right)$. See Methods for the derivation of these results. The functions $\chi(h)$ and $\phi_{i}(h)$ are given as follows. $\chi(h)$ is entropy of $(0,1)$ binary random variables with mean $r(h)=1/(1+e^{-h})$:
$\chi(h)=-r(h)\log r(h)-(1-r(h))\log(1-r(h)).$ (15)
$\phi_{i}(h)$ is given by
$\phi_{i}(h)=-(m_{i,t-1}h-\psi(h)),$ (16)
@@
Here the input $h$ is a Gaussian random variable with mean $g_{i,t,s}$ and variance $\Delta_{i,t,s}$ (s=t,t1s=t,t-1), which can be computed as $h=g_{i,t,s}+z\sqrt{\Delta_{i,t,s}}$ using a standardized Gaussian random variable $z$. $g_{i,t,s}$ and $\Delta_{i,t,s}$ are computed using the mean-field activation rate at the previous step $m_{i,t-1}$ as
@@
The mean-field activation rate $m_{i,t}$ ($i=1,\ldots,N,t=1,\ldots,T$) can be recursively computed using
$m_{i,t}\approx\int\mathcal{D}_{x}\,r(g_{i,t,t-1}+x\sqrt{\Delta_{i,t,t-1}}),$ (19)
@@
We also note that under the steady-state assumption, the mean-field approximation can be expressed using the stationary parameters $m_{i}$, $g_{i}$, and $\Delta_{i}$ as (See Supporting Information 3 text):

### v3

@@ Abstract
Neuronal ensemble activity, including coordinated and oscillatory patterns, exhibits hallmarks of nonequilibrium systems with time-asymmetric trajectories to maintain their organization. However, assessing time asymmetry from neuronal spiking activity remains challenging. The kinetic Ising model provides a framework for studying the causal, nonequilibrium dynamics in spiking recurrent neural networks. Recent theoretical advances in this model have enabled time-asymmetry estimation from large-scale steady-state data. Yet, neuronal activity often exhibits time-varying firing rates and coupling strengths, violating the steady-state assumption. To overcome this limitation, we developed a state-space kinetic Ising model that accounts for nonstationary and nonequilibrium properties of neural systems. This approach incorporates a mean-field method for estimating time-varying entropy flow, a key measure for maintaining the system’s organization through dissipation. Applying this method to mouse visual cortex data revealed greater variability in causal couplings during task engagement despite reduced neuronal activity with increased sparsity. Moreover, higher-performing mice exhibited increased coupling-related entropy flow per spike during task engagement, suggesting more efficient computation in the higher-performing mice. These findings underscore the model’s utility in uncovering intricate asymmetric causal dynamics in neuronal ensembles and linking them to behavior through the thermodynamic underpinnings of neural computation.
@@ Introduction
Neural systems are no exception. In animals engaged in behavioral and cognitive tasks, the dynamics of neuronal population activity exhibit hallmarks of nonequilibrium systems. Notable examples include the rotational activity of M1 neurons during motor execution tasks [16, 17] and the sequential patterns observed in hippocampal neurons, including their replay, during navigation and sleep [18, 19, 20]. Since the original proposal of cell assembly and its phase sequences by Donald O. Hebb [21], coordinated sequential patterns have been thought fundamental for memory consolidation and retrieval [22, 23, 24, 25]. Recently, studies on fMRI or ECoG suggested that increased time-asymmetry in neural signals, quantified by steady-state entropy production [7, 26, 27], could serve as a signature of consciousness [28, 29, 30, 31] or reflect the cognitive load demanded by tasks [32]. For instance, entropy production measured from ECoG signals of non-human primates is diminished during sleep and certain types of anesthesia compared to wakefulness [28, 31], indicating that the awake state includes more directed temporal patterns. However, assessing entropy production directly from neuronal spiking activities remains challenging. Further complicating this issue, neural signals exhibit nonstationary dynamics, including transient or oscillatory behavior, which hinders the use of steady-state entropy production metrics.
@@
The kinetic Ising model is a prototypical model of recurrent neural networks [33, 34]. It extends the equilibrium Ising model, which has been successfully applied to empirical spiking data to elucidate the thermodynamic and associative-memory properties of neural systems [35, 36]. In the kinetic Ising system, neurons are causally driven by the past states of self and other neurons, as well as a force representing the intrinsic excitability of the neurons and/or an influence of unobserved concurrent signals. When neurons receive steady inputs and their causal couplings are asymmetric, the system does not relax to an equilibrium state. Instead, it exhibits steady-state nonequilibrium dynamics characterized by non-zero entropy production. Recent theoretical studies on steady-state entropy production have elucidated its behavior in relation to distinct phases of the Ising system, including critical phase transitions [37]. Mean-field theories have been developed for kinetic Ising systems [38, 39, 40, 41, 42, 43], enabling the estimation of steady-state entropy production from large-scale spike sequences [43]. However, neuronal activity exhibits dynamical changes not only in firing rates but also in the strength of their interactions, both of which violate the steady-state assumptions.
@@
To account for the nonstationary dynamics of neural systems, the state-space method [44, 45] has been applied to the Ising system [46, 47, 48, 49, 50]. In these approaches, Bayesian filtering and smoothing algorithms have been developed to estimate time-dependent parameters of the Ising model, along with an EM algorithm [51, 52] to optimize various hyperparameters. These models have enabled researchers to trace time-varying neuronal interactions while neurons’ internal parameters change dynamically, absorbing the effect from unobserved concurrent signals. Additionally, it has elucidated the thermodynamic quantities of neural systems (e.g., free energy and specific heat) in a time-dependent manner, in relation to the behavioral paradigms of tasks [48]. Nevertheless, these methods assume an equilibrium Ising model with symmetric couplings, which limits their ability to assess the nonequilibrium properties of observed neural activities.
@@
In this study, we develop the state-space kinetic Ising model to account for the nonstationary and nonequilibrium properties of neural activities. We also construct a mean-field method for estimating time-varying entropy flow, an essential component of entropy production that quantifies the dissipation of entropy, from spiking activities of neural ensembles. Given this method, we hypothesize that entropy flow, as estimated with the kinetic Ising framework, reflects the capacity of neural populations to perform meaningful computation under energetic and behavioral-time constraints [53]. Specifically, we expect that high-performing animals would exhibit greater entropy flow per spike, consistent with efficient coding.
@@
Application of the methods to mouse V1 neurons revealed behavior-dependent changes in entropy flow. From the analysis of 37 mice, we found that while spike rates of the populations are lower on average and exhibited sparser distributions when mice actively engaged in tasks than in the passive condition, active engagement significantly enhanced the variability of the neuronal couplings, which contributed to increasing entropy flow. Further, higher-performing mice exhibited stronger entropy flow per spike in active engagement than in the passive condition. We corroborated contributions of couplings to this tendency using trial-shuffled data that excluded influences of firing rate dynamics and sampling errors in estimating neuronal couplings. Thus, the method enabled us to reveal contributions of behavior-related neuronal couplings to the causal activities in sparsely active neuronal populations, while isolating firing rate dynamics. These results imply economical representations of stimuli by time-asymmetric causal activity in competent mice.
@@
This paper is organized as follows. In Results, we first introduce the state-space kinetic Ising model and its estimation method. Next, we introduce the mean-field method for estimating entropy flow. We validate these methods through simulations and then apply them to mouse V1 data. Finally, we relate our findings to previous studies and discuss their implications for efficient information coding in neural populations.
@@ Results
In neurophysiological experiments, the experimentalists simultaneously record the activity of multiple neurons while animals are exposed to a stimulus or perform a task, and repeat the recordings multiple times under the same experimental conditions. We analyze the quasi-simultaneous activity of neurons using binarized spike sequences. For this goal, we convert the simultaneous sequences of spike timings of $N$ neurons into sequences of binary patterns by binning them with a bin width of $\Delta$ [ms]. We assign a value of $1$ if there is one or more spikes in a bin and 0 otherwise. We assume that there are $T+1$ bins for each trial, with an initial bin being the 0-th bin and $L$ trials in total. Below, we treat the bins as discrete time steps and refer to the $t$-th bin as time $t$. We let $x^{l}_{i,t}=\{0,1\}$ be a binary variable of the $i$-th neuron at time $t$ in the $l$-th trial ($i=1,\ldots,N$, $t=0,\ldots,T$, $l=1,\ldots,L$). We collectively denote the binary patterns of simultaneously recorded neurons at time $t$ in the $l$-th trial using a vector, $\mathbf{x}^{l}_{t}=(x^{l}_{1,t},\ldots,x^{l}_{i,t},\ldots,x^{l}_{N,t})$. Further, we denote the patterns at time $t$ from all trials by $\mathbf{x}_{t}=(\mathbf{x}^{1}_{t},\ldots,\mathbf{x}^{l}_{t},\ldots,\mathbf{x}^{L}_{t})$ and denote all the patterns up to time $t$ by $\mathbf{x}_{0:t}$.
@@
where $\theta_{i,t}$ is a time-dependent (external) field parameter that determines the bias for inputs to the $i$-th neuron at time $t$ and $\theta_{ij,t}$ is a time-dependent coupling parameter from the $j$-th neuron to the $i$-th neuron. These parameters are collectively denoted as $\boldsymbol{\theta}_{t}=(\boldsymbol{\theta}^{1}_{t},\ldots,\boldsymbol{\theta}^{i}_{t},\ldots,\boldsymbol{\theta}^{N}_{t})$ and $\boldsymbol{\theta}^{i}_{t}=(\theta_{i,t},\theta_{i1,t},\ldots\theta_{ij,t},\ldots,{\theta}_{iN,t})$. $\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$ is the log normalization term defined as
@@
We also specify $p(\mathbf{x}_{0})$, a probability mass function of the binary patterns at $t=0$, which we assume $p(\mathbf{x}_{0})=\prod_{i=1}^{N}\prod_{l=1}^{L}p(x^{l}_{i,0})$, where $p(x^{l}_{i,0})=0.5$ for data generation.
@@
$p(\boldsymbol{\theta}_{1:T}|\mathbf{w})$ $=\prod_{i=1}^{N}\left[p(\boldsymbol{\theta}_{1}^{i}|\boldsymbol{\mu}^{i},\boldsymbol{\Sigma}^{i})\prod_{t=2}^{T}p(\boldsymbol{\theta}_{t}^{i}|\boldsymbol{\theta}_{t-1}^{i},\mathbf{Q}^{i})\right],$ (3)
where $\mathbf{w}$ denotes the collection of the hyperparameters: $\mathbf{w}=[\boldsymbol{\mu}^{1},\ldots,\boldsymbol{\mu}^{N},\boldsymbol{\Sigma}^{1},\ldots,\boldsymbol{\Sigma}^{N},\mathbf{Q}^{1},\ldots,\mathbf{Q}^{N}]$. Namely, we assume independence of the parameters of a neuron from those of the other neurons, which significantly reduces computational costs. The transition of the $i$-th neuron follows the linear Gaussian models:
$p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1},\mathbf{Q}^{i})$
$=\frac{1}{\sqrt{|2\pi\mathbf{Q}^{i}|}}\exp\left[\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})\right],$ (4)
while the initial density $p(\boldsymbol{\theta}_{1}^{i}|\boldsymbol{\mu}^{i},\boldsymbol{\Sigma}^{i})$ is given by the Gaussian distribution with mean $\boldsymbol{\mu^{i}}$ and covariance $\boldsymbol{\Sigma^{i}}$.
Model fitting and inference
@@
while optimizing the hyperparameters $\mathbf{w}$ under the principle of maximizing marginal likelihood:
@@
The Expectation-Maximization (EM) algorithm [54] offers a way to construct the approximate posterior with optimized hyperparameters by alternately constructing the approximate posterior density while fixing the hyperparameters (E-step) and optimizing the hyperparameters while fixing the approximate posterior (M-step). The construction of the approximate posterior density at the E-step is performed by sequentially applying Bayes algorithms in a forward and backward manner, where we approximate the posteriors by Gaussian distributions using Laplace’s method. Thus, the method yields the mean and variance of the approximated Gaussian posterior at time $t$, which are denoted as $\boldsymbol{\theta}_{t|T}$ and $\mathbf{W}_{t|T}$, respectively. See Methods and Supplementary Note 1 for the details of the algorithm.
Entropy flow
Using the inferred parameters $\boldsymbol{\theta}_{1:T}$ of the kinetic Ising model from spike data, we estimate entropy flow (also known as bath entropy change) at each time step. The entropy flow at time $t$ is defined as:
$\sigma_{t}^{\rm flow}=\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}p(\mathbf{x}_{t},\mathbf{x}_{t-1})\log\frac{p(\mathbf{x}_{t}{|}\mathbf{x}_{t-1})}{p(\mathbf{x}_{t-1}{|}\mathbf{x}_{t})},$ (7)
where $p(\mathbf{x}_{t-1}{|}\mathbf{x}_{t})$ represents the probability of observing time-reversed processes generated under the forward model. Because we use the natural logarithm, we report entropy flow in units of nats. Eq. 7 is related to the entropy production $\sigma_{t}$ at time $t$ as follows [7, 9, 26, 27]:
$\sigma_{t}$ $=\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}p(\mathbf{x}_{t},\mathbf{x}_{t-1})\log{\frac{p(\mathbf{x}_{t}{|}\mathbf{x}_{t-1})p_{t-1}(\mathbf{x}_{t-1})}{p(\mathbf{x}_{t-1}{|}\mathbf{x}_{t})p_{t}(\mathbf{x}_{t})}}$
@@
The entropy production is non-negative: $\sigma_{t}\geq 0$. Thus, the positive entropy flow allows a decrease in the system’s entropy: namely, the system can be more structured or organized when the entropy flow is positive. Since it is challenging to estimate the system’s entropy or its change, here we estimate the entropy flow, which provides the lower bound of the entropy change: $S_{t}-S_{t-1}\geq-\sigma_{t}^{\rm flow}$.
Similarly, since the total entropy production across all time steps ${\sigma}_{1:T}$ is given as
@@
In this study, we refer to Eq. 7 as entropy flow because it is related to heat flow to reservoirs (thermal bath) and the entropy change of the reservoirs in thermodynamics [53]. We note that Eq. 7 differs from the entropy flow defined in [55, 56], which was obtained by the decomposition of the dissipation function [27] as an alternative to entropy production. See [27, 57] for their distinct definitions and decompositions for the case of discrete-time systems.
@@
$\sigma_{t}^{\rm flow}$ $=\sum_{i}\theta_{i,t}\left(E_{\mathbf{x}_{t}}x_{i,t}-E_{\mathbf{x}_{t-1}}x_{i,t-1}\right)$
$\phantom{=}+\sum_{i,j}\theta_{ij,t}E_{\mathbf{x}_{t},\mathbf{x}_{t-1}}\left(x_{i,t}x_{j,t-1}-x_{i,t-1}x_{j,t}\right)$
$\phantom{=}-\sum_{i}\left(E_{\mathbf{x}_{t-1}}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{t-1})-E_{\mathbf{x}_{t}}\psi(\boldsymbol{\theta}^{i}_{t-1},\mathbf{x}_{t})\right),$ (11)
where $E_{\mathbf{x}_{t}}$ and $E_{\mathbf{x}_{t},\mathbf{x}_{t-1}}$ represents expectation by $p(\mathbf{x}_{t})$ and $p(\mathbf{x}_{t},\mathbf{x}_{t-1})$, respectively.
Mean-field estimation of entropy flow
Entropy flow (Eq. 7) requires the expectation by the joint density $p(\mathbf{x}_{t},\mathbf{x}_{t-1})$, which is computationally expensive for large systems. While the mean-field methods for the kinetic Ising model [38, 39, 40, 41, 42, 43] were employed to estimate steady-state entropy flow [43], the mean-field method for estimating time-varying entropy flow remains unexplored. Here, we develop the mean-field method for estimating dynamic entropy flow.
The entropy flow $\sigma_{t}^{\rm flow}$ can be decomposed into the forward and reverse components,
@@
where $\sigma_{t}^{\rm forward}$ and $\sigma_{t}^{\rm backward}$ denote the conditional entropies of the forward and time-reversed conditional distributions, respectively. The proposed mean-field method estimates the entropy flow by approximating the forward and time-reversed conditional entropies using the Gaussian integral:
@@
${\sigma}^{\rm backward}_{t}$ $\approx\sum_{i=1}^{N}\int\mathcal{D}_{z}\,\phi_{i,t}(g_{i,t,t}+z\sqrt{\Delta_{i,t,t}}),$ (14)
where $\mathcal{D}_{z}=\frac{\mathrm{d}z}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}z^{2}\right)$. See Methods and Supplementary Note 2 for the derivation of these results. The functions $\chi(h)$ and $\phi_{i,t}(h)$ are given as follows. $\chi(h)$ is entropy of $(0,1)$ binary random variables with mean $r(h)=1/(1+e^{-h})$:
$\chi(h)=-r(h)h+\psi(h),$ (15)
where we redefined the log normalization function $\psi$ as a function of $h$: $\psi(h)=\log(1+e^{h})$. $\phi_{i,t}(h)$ is given by
$\phi_{i,t}(h)=-m_{i,t-1}h+\psi(h),$ (16)
@@
Here, the input $h=g_{i,t,s}+z\sqrt{\Delta_{i,t,s}}$ is a Gaussian random variable with mean $g_{i,t,s}$ and variance $\Delta_{i,t,s}$ (s=t,t1s=t,t-1), where $z$ denotes a standardized Gaussian random variable. $g_{i,t,s}$ and $\Delta_{i,t,s}$ are computed using the mean-field activation rate at time $s$, $m_{i,s}$, as
@@
The mean-field activation rate $m_{i,t}$ can be recursively computed using
$m_{i,t}\approx\int\mathcal{D}_{z}\,r(g_{i,t,t-1}+z\sqrt{\Delta_{i,t,t-1}}),$ (19)
@@
[…]
We also note that under the steady-state assumption, the mean-field approximation can be expressed using the stationary parameters $m_{i}$, $g_{i}$, and $\Delta_{i}$ as (See Supplementary Note 3):